40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 27 03:56:52 2015, ericq, Update, LSC, DRFPMI work 
    Reply  Thu Oct 29 03:25:49 2015, ericq, Update, LSC, DRFPMI work DRFPMI.pdfDRMIarms.pdf
       Reply  Tue Nov 3 03:12:46 2015, ericq, Update, LSC, DRFPMI work 
Message ID: 11718     Entry time: Tue Oct 27 03:56:52 2015     Reply to this: 11722
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI work 

A handful of DRFPMI locks tonight, longest one was ~7 minutes. 

EPICS/network latency has been a huge pain tonight. The locking script may hang between commands at an unstable place, or fail to execute commands altogether because it can't find the EPICS channel. This prevented or broke a number of locks. 

I made some CARM OLG and crossover measurements, and found the AO gain for the right crossover freq (~100Hz) to be ~8dB different than what's in the PRFPMI script, which is weird. Right now, the CARM bandwidth / ability to turn on boosts is limited by the gain peaking in the IMC CLG due to the high-ish PC/PZT crossover frequency we're using. 

Gautam turned on some sensing excitations during the last couple of locks, but they weren't on for very long before the lock loss. Hopefully I can pull out at least some angles from the data. 

I'm also more convinced that the PRC angular FF needs retuning; there is more residual motion on the cameras than I'm used to seeing. I've taken more data that I'll use to recalculate a wiener filter tomorrow. 

The PMC, ALSX beat and ITMX oplev all needed a reasonable pitch realignment tonight. 

ELOG V3.1.3-